OpenStudy (anonymous):

use the communtive property to rewrite the expresdion 5x + (-3y)

OpenStudy (usukidoll):

don't you mean commutative?

OpenStudy (usukidoll):

commutative law A+B = B +A Let A = 5x and B = -3y Using the commutative law, we have -3y+5x

OpenStudy (usukidoll):

UGH! dang.. that's what I forgot to write on my proof! :O

OpenStudy (anonymous):

it is equal to (-3y)+5x

OpenStudy (usukidoll):

I just wrote a basic proof on here it is -3y+5x

OpenStudy (anonymous):

yeah!! I thought i woulnt need to show proof!

OpenStudy (anonymous):

thank you both !